OpenStudy (anonymous):

Two cars start from the same point at the same time and travel in opposite directions. The slow car travels 56 miles per hr and the fast car travels at 60 miles per hour. In how many hours will they be 464 miles apart?

OpenStudy (cwrw238):

combined speed = 116 mph time in hours to travel 464 miles = 464/116

OpenStudy (anonymous):

I don't understand how you got that answer. I am sorry

OpenStudy (cwrw238):

156 + 60 = 116 mph the speed they are travelling apart is 116 mph speed =distance / time 116 = 464 / t 116t = 464 t = 464/116 = 4 hours